OpenStudy (anonymous):

7x-7y=42 graph the question

OpenStudy (anonymous):

let x=0 then y=-6 => (0,-6) let y=0 then x=6 => (6,0) graph to follow

OpenStudy (anonymous):

how about the other question like y= 1-3x can u graph that problem for me

OpenStudy (anonymous):

Sorry mandii, yours was the last question I answered at that sitting, and I left before you posted your second question. I'll answer it in case you didn't get a good answer after from someone else I left. :{ y-int: (0, 1) from y=mx+b form x-int: let y=0 and solve for x: 0=1-3x 3x=1 x=1/3 => (1/3, 0) graph attached below